Differentiate the following functions:

y = (t^3 - 5)^1/a


1
Expert's answer
2021-05-11T07:10:06-0400

Given function is-


"y=(t^3-5)^{\frac{1}{a}}"


Differentiate with respect to "x-"


"\dfrac{dy}{dt}=\dfrac{1}{a}(t^3-5)^{\frac{1}{a}-1}\times 3t^2\\[9pt]y'(t)=\dfrac{3t^2}{a}(t^3-5)^{\frac{1-a}{a}}"